OpenStudy (anonymous):

Find the extreme values of each function on the interval!! 10 points!!? ³√(x) ; -2 ≤ x ≤ 27 answer: C) max= (27,3) and min = (0,0) This is the last problem I have and I cannot figure out how to get this answer! Can someone show me please??? Thanks in advance!!

OpenStudy (anonymous):

Sorry the question says "find the ABSOLUTE extreme values of each function on the interval"

OpenStudy (mosaic):

f(x) = ³√(x) f'(x) = 1/3x^(-2/3) = 1/(3x^2/3) which is positive for all values of x and is 0 at x = 0. Therefore, f(x) is an increasing function. In any interval, it will attain its maximum at the right endpoint of that interval because the function always increases as x increases.

OpenStudy (mosaic):

In the interval [-2,27], f(x) will attain its maximum when x = 27 f(27) = (27)^(1/3) = 3. The max point is (27,3). When x = 0, f(x) = 0. When we take the absolute value, all the negative f(x) will become positive and the lowest or the minimum will be 0. Therefore, absolute minimum occurs at x = 0 and the point is (0,0).